OpenStudy (anonymous):

>>NEED FAST HELP, WILL FAN AND MEDAL<< y= 3/4x+10 What is x?

OpenStudy (anonymous):

@imyoursquirtle

OpenStudy (anonymous):

x usually means the input value, and y means the output value of an equation ^_^

OpenStudy (anonymous):

@DemolisionWolf I know what it is, but I mean solve for x.

OpenStudy (anonymous):

make the equation equal to 'x' not 'y'?

OpenStudy (anonymous):

@DemolisionWolf

OpenStudy (anonymous):

oooohh i see. the x intercept means y=0 y= 3/4x+10 becomes 0= 3/4x+10 step 1) get the 'x term' alone , can you do this?

OpenStudy (anonymous):

@DemolisionWolf Err, no...

OpenStudy (anonymous):

0= 3/4x+10 hint* subtract 10 from both sides

OpenStudy (anonymous):

-10 and 0??? @DemolisionWolf

OpenStudy (anonymous):

0= 3/4x+10 like this: 0 - 10= 3/4x+10 -10 now simplify it ^_^

OpenStudy (anonymous):

10 = 3/4 ??!!?

OpenStudy (anonymous):

close! let just look at each side by its self 0-10 = -10 and 3/4x+10 -10 = 3/4x so it is -10 = 3/4x now we need to get the x alone completely, so times both sides by 4

OpenStudy (anonymous):

-10 times 4 = -40 3/4 times 4 = 3 @DemolisionWolf

OpenStudy (anonymous):

excellent ^_^

OpenStudy (anonymous):

-40 = 3x is what we now have now divide both sides by 3

OpenStudy (anonymous):

-40 = -13 3 = 1

OpenStudy (anonymous):

do -40/3 =? 3x/3=?

OpenStudy (anonymous):

I did. -13 & 1 were my answers.

OpenStudy (anonymous):

oh, you got it, ya -13.33 and 1 are correct, but remember the 'x' with the 1, so it's 1x

OpenStudy (anonymous):

Okay, so whats the answer out of these choices? A. -13.3,0 B. 0,-13.3 C. 0,10 D. 10,0

OpenStudy (anonymous):

alright, so remember way back at the start, when I said that an x-intercept means y=0? so we know the that the answer has to look like this: ( # , 0) and we just finished solving and got x=-13.3 remember?

OpenStudy (anonymous):

So, it's A?

OpenStudy (anonymous):

yep ^_^